OpenStudy (anonymous):

find an equation of the sphere with center (2,-6,4) and radius 5. Describe its intersection with each of the coordinates planes.

OpenStudy (anonymous):

so i know how to do the equation, but how do i do the intersection?

OpenStudy (anonymous):

\[(x-2)^2 + (y+6)^2 + (z-4)^2 = 25\]

hero (hero):

What course are you taking?

OpenStudy (anonymous):

calc 3

OpenStudy (anonymous):

the description of the intersection is an equation of a circle i believe based on the example prior to this problem

OpenStudy (amistre64):



OpenStudy (amistre64):

you know the distance you are from the plane; you know the Radius of the sphere, so we can calculate the radius of the circle in the plane

OpenStudy (amistre64):

the plane you pick is just dropping that variable from the picture to get your center in that plane: (1,2,3) centers in no-x plane at (2,3) the no-y plane at (1,3) and the no-z plane at (1,2)

OpenStudy (amistre64):

the distance away from the plane is the one that you dropped :)

OpenStudy (anonymous):

okay so the xy plane would be (x-2)^2 + *y+6)^2 = 9

OpenStudy (anonymous):

The sphere does not intersect the y=0 coordinate plane.

OpenStudy (anonymous):

wait it would be = 16?

OpenStudy (amistre64):

(2,-6,4) and radius 5 xy: no-z: center at (2,-6) with a distance of 4 radius in no-z = sqrt(5^2-4^2)

OpenStudy (anonymous):

ohhh okay okay

OpenStudy (anonymous):

wait still confused

OpenStudy (anonymous):

so what would be the equation for yz?

hero (hero):

How do I unsubscribe from this?

OpenStudy (amistre64):

yz: no-x plane; centers on (y,z) with a radius of: sqrt(5^2 - x^2)

OpenStudy (anonymous):

so 25-4? root 21?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

and r^2 = 21 then in your equation of circle stuff

OpenStudy (anonymous):

but how would i do xz?

OpenStudy (anonymous):

r = root(25-36)

OpenStudy (amistre64):

well, it centers in the no-y so (x,z); its radius is then sqrt(5^2-y^2)

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

but its a negative number

OpenStudy (amistre64):

it is isnt it :) tell me; if your stretching to reach 5 and your 6 away, do you hit it?

OpenStudy (anonymous):

yeah bc r =6?

OpenStudy (amistre64):



OpenStudy (anonymous):

oh so it never intersects?

OpenStudy (amistre64):

its negative meaning there is no real radius in that plane so we aint even touching it

OpenStudy (anonymous):

so there wouldn't be an equation for xz

OpenStudy (amistre64):

correct; the xz plane has no intersection with the sphere

OpenStudy (anonymous):

awesome! thank you so much!

OpenStudy (amistre64):

yep